OpenStudy (anonymous):

how i can prove 222221 is composite

OpenStudy (vishweshshrimali5):

222221 = 359 * 619 Since it has other factors also except 1 and the number itself, thus it is a composite number

ganeshie8 (ganeshie8):

@ikram002p try this when free...

ganeshie8 (ganeshie8):

one failed attempt : 222221 is NOT a prime if we could show that : \[\large a^{222221 } \not \equiv a \mod 222221 \] I have tried to reduce using a=2, 3... but its not simplifying : \[\large 2^{222221 } \not \equiv 2 \mod 222221 \]

OpenStudy (ikram002p):

mm ganesh im not sure the inverse of frmats work mmm how ever it has factorz mmm

OpenStudy (ikram002p):

fermats set if p is prime then a^p= a mod p , but that does not means if there exist n such that a^n=a mod n for any n then n should be prime :o

ganeshie8 (ganeshie8):

we can use contrapositive right ? \(p\) is NOT a prime number if : \[\large a^p \not \equiv a \mod p\] for some \(a\)

OpenStudy (ikram002p):

yeah that works xD

ganeshie8 (ganeshie8):

we're not proving \(p\) is a prime number, we're trying to prove \(p\) is NOT a prime number

ganeshie8 (ganeshie8):

I have tried a=2, 3 and gave up on this already... >.<

OpenStudy (ikram002p):

well lets try wilson :o

OpenStudy (ikram002p):

(p-1) ! =-1 mod p

ganeshie8 (ganeshie8):

Oh never used wilson for prime testing before xD

OpenStudy (ikram002p):

i think mmm we only need to check if (222220 ) ! = -1 mod 222221

OpenStudy (ikram002p):

hehe but that sounds like more factorization xD

ganeshie8 (ganeshie8):

lets go back to FLT

OpenStudy (ikram002p):

yeah :P findint counter example

OpenStudy (ikram002p):

well i bet 359, 619 wont work , right ?

OpenStudy (ikram002p):

4 also dnt work :O

ganeshie8 (ganeshie8):

4 and 2 are same

OpenStudy (ikram002p):

3 also :o

OpenStudy (ikram002p):

well it does not work for 2,3,4,5,6 :P thats enough right ?

ganeshie8 (ganeshie8):

yeah thats more than enough to give up :P

OpenStudy (ikram002p):

lolz ok u already set it up :P

OpenStudy (ikram002p):

then \(\huge\cal\color{pink}{^❤ _❤❤done_❤❤^❤_❤} \)

OpenStudy (crashonce):

well 222221 is divisible by 359 so you can say it is composite